# Dual-Head Knowledge Distillation: Enhancing Logits Utilization with an Auxiliary Head

### v1

@@
Abstract
Traditional knowledge distillation focuses on aligning the student’s predicted probabilities with both ground-truth labels and the teacher’s predicted probabilities. However, the transition to predicted probabilities from logits would obscure certain indispensable information. To address this issue, it is intuitive to additionally introduce a logit-level loss function as a supplement to the widely used probability-level loss function, for exploiting the latent information of logits. Unfortunately, we empirically find that the amalgamation of the newly introduced logit-level loss and the previous probability-level loss will lead to performance degeneration, even trailing behind the performance of employing either loss in isolation. We attribute this phenomenon to the collapse of the classification head, which is verified by our theoretical analysis based on the neural collapse theory. Specifically, the gradients of the two loss functions exhibit contradictions in the linear classifier yet display no such conflict within the backbone. Drawing from the theoretical analysis, we propose a novel method called dual-head knowledge distillation, which partitions the linear classifier into two classification heads responsible for different losses, thereby preserving the beneficial effects of both losses on the backbone while eliminating adverse influences on the classification head. Extensive experiments validate that our method can effectively exploit the information inside the logits and achieve superior performance against state-of-the-art counterparts.
1 Introduction
@@
Figure 1: The reason for introducing the BinaryKL loss and the incompatibility of the CE loss and the BinaryKL loss. Figure 1(a) shows that two different vectors may become the same through the softmax function, which means some information would be lost during the transformation process carried out by the softmax function. Figure 1(b) shows four settings we evaluate: (1) only cross-entropy (CE) loss; (2) only the BinaryKL loss; (3) CE + BinaryKL; (4) Our proposed Dual-Head Knowledge Distillation (DHKD). The red cross over the third setting means that the student model trained under this setting will collapse. The performance sorting of four settings is (4) $>$ (1) $>$ (2) $\gg$ (3), as shown in Figure 2.
@@
2 Related Work
2.1 Knowledge Distillation
@@
Feature-based methods focus on distilling knowledge from intermediate feature layers. FitNet (Romero et al., 2015) is the first approach to distill knowledge from intermediate features by measuring the distance between feature maps. RKD (Park et al., 2019) utilizes the relations among instances to guide the training process of the student model. CRD (Tian et al., 2019) incorporates contrastive learning into knowledge distillation. OFD (Heo et al., 2019) contains a new distance function to distill significant information between the teacher and student using marginal ReLU. ReviewKD (Chen et al., 2021) proposes a review mechanism that uses multiple layers in the teacher to supervise one layer in the student. Other papers (Passalis & Tefas, 2018; Kim et al., 2018; Koratana et al., 2019; Li, 2022; Liu et al., 2023; Wang et al., 2023; Roy Miles & Deng, 2024; Miles & Mikolajczyk, 2024) enforce various criteria based on features. Most feature-based methods can attain superior performance, yet involving considerably high computational and storage costs.
Logit-based methods mainly concentrate on distilling knowledge from logits and softmax scores after logits. DML (Zhang et al., 2018b) introduces a mutual learning method to train both teachers and students simultaneously. DKD (Zhao et al., 2022) proposes a novel logit-based method to reformulate the classical KD loss into two parts and achieves state-of-the-art performance by adjusting weights for these two parts. DIST (Huang et al., 2022) relaxes the exact matching in previous KL divergence loss with a correlation-based loss and performs better when the discrepancy between the teacher and the student is large. TTM (Zheng & Yang, 2024) drops the temperature scaling on the student side, which causes an inherent Renyi entropy term as an extra regularization term in the loss function. Although logit-based methods require fewer computational and storage resources, they experience a performance gap compared with feature-based methods.
2.2 Neural Collapse
@@
Although the phenomenon is intuitive, its reason has not been entirely understood, which inspires several lines of theoretical work on it. Papyan et al. (2020) prove that if the features satisfy neural collapse, the optimal classifier vectors under the MSE loss will also converge to neural collapse. Some studies turn to a simplified model that only considers the last-layer features and the classifier as independent variables, and they prove that neural collapse emerges under the CE loss with proper constraints or regularization (Fang et al., 2021; Ji et al., 2022; Lu & Steinerberger, 2022; Weinan E, 2022; Zhu et al., 2021). Yang et al. (2022) also use such a simplified model and decompose the gradient into two parts to help build a better classifier for class-imbalanced learning, which is the primary technical reference for the theoretical analysis section of this article.
2.3 Decoupled Heads
@@
3 Dual-Head Knowledge Distillation
@@
3.1 Incompatibility between CE and BinaryKL
Figure 2: The test accuracy and test loss curves of the student models during the training phase. We set resnet56 as the teacher and resnet20 as the student on the CIFAR-100 dataset.
@@ 3.1 Incompatibility between CE and BinaryKL
$\mathcal{L}_{\mathrm{BinaryKL}}=\tau^{2}\sum\limits_{i=1}^{B}\sum\limits_{k=1}^{K}\mathcal{KL}\left(\left[\sigma\left(z_{i,k}^{\mathcal{T}}/\tau\right),1-\sigma\left(z_{i,k}^{\mathcal{T}}/\tau\right)\right]\|\left[\sigma\left(z_{i,k}^{\mathcal{S}}/\tau\right),1-\sigma\left(z_{i,k}^{\mathcal{S}}/\tau\right)\right]\right),$ (1)
@@
As shown in Figures 1 and 2, although the BinaryKL loss performs well alone, it is incompatible with the CE loss. The student model can achieve descent performance by training with either the CE loss or the BinaryKL loss separately, but the performance will be degraded severely when we combine them as follows:
$\mathcal{L}_{\mathrm{overall}}=\mathcal{L}_{\mathrm{CE}}+\alpha\mathcal{L}_{\mathrm{BinaryKL}}.$ (2)
@@
(a) A simplex equiangular tight frame
(b) Gradient directions w.r.t. $\bm{w}$
(c) Gradient directions w.r.t. $\bm{h}$
Figure 3: Gradient analysis based on the neural collapse theory. (a) An illustration of a simplex ETF when $d=3$ and $K=4$. The “+” and “” with different colors refer to features and classifier vectors of different classes, respectively. (b) Gradient directions about a certain $\bm{h}$ (belongs to the 1-st class) w.r.t. all $\bm{w}_{i},i\in{1,2,3}$. (c) Gradient directions w.r.t. an $\bm{h}$ (belongs to the 1-st class).
3.2 Theoretical Analysis
@@
Papyan et al. (2020) revealed the neural collapse phenomenon, where the last-layer features converge to their within-class means, and the within-class means together with the classifier vectors collapse to the vertices of a simplex equiangular tight frame (ETF) at the terminal phase of training. According to Papyan et al. (2020), all vectors in a simplex ETF have an equal $\ell_{2}$ norm and the same pairwise angle. An illustration of a simplex ETF is shown in Figure 3. The neural collapse phenomenon (Papyan et al., 2020) can be characterized by four manifestations: (1) the variability of the last-layer features in the same class collapses to zero; (2) the class means of different classes’ last-layer features converge to a simplex ETF; (3) the linear classifier vectors collapse to the class means; (4) the task of predicting for classification can be simplified into finding out the nearest class center of the last-layer feature. The detailed definition of a simplex ETF and the features of the neural collapse phenomenon can be found in Appendix B.1.
@@
$\frac{\partial\mathcal{L}_{\mathrm{overall}}}{\partial\bm{w}_{k}}=-(\bm{F}_{\bm{w}\mathrm{\text{-}pull}}^{\mathrm{CE}}+\alpha\bm{F}_{\bm{w}\mathrm{\text{-}pull}}^{\mathrm{BinaryKL}})-(\bm{F}_{\bm{w}\mathrm{\text{-}push}}^{\mathrm{CE}}+\alpha\bm{F}_{\bm{w}\mathrm{\text{-}push}}^{\mathrm{BinaryKL}}),$ (3)
@@
$\bm{F}_{\bm{w}\mathrm{\text{-}pull}}^{\mathrm{CE}}=\sum\nolimits_{i=1}^{n_{k}}(1-p_{k}(\bm{h}^{\mathcal{S}}_{k,i}))\bm{h}^{\mathcal{S}}_{k,i},$ $\bm{F}_{\bm{w}\mathrm{\text{-}pull}}^{\mathrm{BinaryKL}}=\tau\sum\nolimits_{i=1}^{n_{k}}(q_{k}(\bm{h}^{\mathcal{T}}_{k,i})-q_{k}(\bm{h}^{\mathcal{S}}_{k,i}))\bm{h}^{\mathcal{S}}_{k,i},$
$\bm{F}_{\bm{w}\mathrm{\text{-}push}}^{\mathrm{CE}}=-\sum\limits_{k^{\prime}\neq k}^{K}\sum\limits_{j=1}^{n_{k^{\prime}}}p_{k}(\bm{h}^{\mathcal{S}}_{k^{\prime},i})\bm{h}^{\mathcal{S}}_{k^{\prime},j},$ $\bm{F}_{\bm{w}\mathrm{\text{-}push}}^{\mathrm{BinaryKL}}=-\tau\sum\limits_{k^{\prime}\neq k}^{K}\sum\limits_{j=1}^{n_{k^{\prime}}}(q_{k}(\bm{h}^{\mathcal{S}}_{k,i})-q_{k}(\bm{h}^{\mathcal{T}}_{k,i}))\bm{h}^{\mathcal{S}}_{k^{\prime},j}.$
@@
$\frac{\partial\mathcal{L}_{\mathrm{overall}}}{\partial\bm{h}}=-(\bm{F}_{\bm{h}\mathrm{\text{-}pull}}^{\mathrm{CE}}+\alpha\bm{F}_{\bm{h}\mathrm{\text{-}pull}}^{\mathrm{BinaryKL}})-(\bm{F}_{\bm{h}\mathrm{\text{-}push}}^{\mathrm{CE}}+\alpha\bm{F}_{\bm{h}\mathrm{\text{-}push}}^{\mathrm{BinaryKL}}),$ (4)
@@
$\bm{F}_{\bm{h}\mathrm{\text{-}pull}}^{\mathrm{CE}}=(1-p_{c}(\bm{h}^{\mathcal{S}}))\bm{w}^{\mathcal{S}}_{c},\quad$ $\bm{F}_{\bm{h}\mathrm{\text{-}pull}}^{\mathrm{BinaryKL}}=\tau(q_{c}(\bm{h}^{\mathcal{T}})-q_{c}(\bm{h}^{\mathcal{S}}))\bm{w}^{\mathcal{S}}_{c},$
$\bm{F}_{\bm{h}\mathrm{\text{-}push}}^{\mathrm{CE}}=-\sum\limits_{k\neq c}^{K}p_{k}(\bm{h}^{\mathcal{S}})\bm{w}^{\mathcal{S}}_{k},\quad$ $\bm{F}_{\bm{h}\mathrm{\text{-}push}}^{\mathrm{BinaryKL}}=-\tau\sum\limits_{k\neq c}^{K}(q_{k}(\bm{h}^{\mathcal{S}})-q_{k}(\bm{h}^{\mathcal{T}}))\bm{w}^{\mathcal{S}}_{k}.$
@@
3.3 Our Proposed Method

### v2

@@
Abstract.
Traditional knowledge distillation focuses on aligning the student’s predicted probabilities with both ground-truth labels and the teacher’s predicted probabilities. However, the transition to predicted probabilities from logits would obscure certain indispensable information. To address this issue, it is intuitive to additionally introduce a logit-level loss function as a supplement to the widely used probability-level loss function, for exploiting the latent information of logits. Unfortunately, we empirically find that the amalgamation of the newly introduced logit-level loss and the previous probability-level loss will lead to performance degeneration, even trailing behind the performance of employing either loss in isolation. We attribute this phenomenon to the collapse of the classification head, which is verified by our theoretical analysis based on the neural collapse theory. Specifically, the gradients of the two loss functions exhibit contradictions in the linear classifier yet display no such conflict within the backbone. Drawing from the theoretical analysis, we propose a novel method called dual-head knowledge distillation, which partitions the linear classifier into two classification heads responsible for different losses, thereby preserving the beneficial effects of both losses on the backbone while eliminating adverse influences on the classification head. Extensive experiments validate that our method can effectively exploit the information inside the logits and achieve superior performance against state-of-the-art counterparts. Our code is available at: https://github.com/penghui-yang/DHKD.
Knowledge Distillation, Neural Collapse
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3736915isbn: 979-8-4007-1454-2/2025/08ccs: Computing methodologies Neural networksccs: Computing methodologies Supervised learning by classification
1. Introduction
@@
Figure 1. The reason for introducing the BinaryKL loss and the incompatibility of the CE loss and the BinaryKL loss. Figure 1(a) shows that two different vectors may become the same through the softmax function, which means some information would be lost during the transformation process carried out by the softmax function. Figure 1(b) shows four settings we evaluate: (1) only cross-entropy (CE) loss; (2) only the BinaryKL loss; (3) CE + BinaryKL; (4) Our proposed Dual-Head Knowledge Distillation (DHKD). The red cross over the third setting means that the student model trained under this setting will collapse. The performance sorting of four settings is (4) $>$ (1) $>$ (2) $\gg$ (3), as shown in Figure 2.
[…]
@@
2. Related Work
2.1. Knowledge Distillation
@@
Feature-based methods focus on distilling knowledge from intermediate feature layers. FitNet (Romero et al., 2015) is the first approach to distill knowledge from intermediate features by measuring the distance between feature maps. RKD (Park et al., 2019) utilizes the relations among instances to guide the training process of the student model. CRD (Tian et al., 2019) incorporates contrastive learning into knowledge distillation. OFD (Heo et al., 2019) contains a new distance function to distill significant information between the teacher and student using marginal ReLU. ReviewKD (Chen et al., 2021) proposes a review mechanism that uses multiple layers in the teacher to supervise one layer in the student. Other papers (Passalis and Tefas, 2018; Kim et al., 2018; Koratana et al., 2019; Li, 2022; Liu et al., 2023; Wang et al., 2023; Roy Miles and Deng, 2024; Miles and Mikolajczyk, 2024) enforce various criteria based on features. Most feature-based methods can attain superior performance, yet involving considerably high computational and storage costs.
Logit-based methods mainly concentrate on distilling knowledge from logits and softmax scores after logits. DML (Zhang et al., 2018a) introduces a mutual learning method to train both teachers and students simultaneously. DKD (Zhao et al., 2022) proposes a novel logit-based method to reformulate the classical KD loss into two parts and achieves state-of-the-art performance by adjusting weights for these two parts. DIST (Huang et al., 2022) relaxes the exact matching in previous KL divergence loss with a correlation-based loss and performs better when the discrepancy between the teacher and the student is large. TTM (Zheng and Yang, 2024) drops the temperature scaling on the student side, which causes an inherent Renyi entropy term as an extra regularization term in the loss function. Although logit-based methods require fewer computational and storage resources, they experience a performance gap compared with feature-based methods.
2.2. Neural Collapse
@@
Although the phenomenon is intuitive, its reason has not been entirely understood, which inspires several lines of theoretical work on it. Papyan et al. (2020) prove that if the features satisfy neural collapse, the optimal classifier vectors under the MSE loss will also converge to neural collapse. Some studies turn to a simplified model that only considers the last-layer features and the classifier as independent variables, and they prove that neural collapse emerges under the CE loss with proper constraints or regularization (Fang et al., 2021; Ji et al., 2022; Lu and Steinerberger, 2022; Weinan E, 2022; Zhu et al., 2021). Yang et al. (2022) also use such a simplified model and decompose the gradient into two parts to help build a better classifier for class-imbalanced learning, which is the primary technical reference for the theoretical analysis section of this article.
2.3. Decoupled Heads
@@
3. Dual-Head Knowledge Distillation
@@
3.1. Incompatibility between CE and BinaryKL
Figure 2. The test accuracy and test loss curves of the student models during the training phase. We set resnet56 as the teacher and resnet20 as the student on the CIFAR-100 dataset.
[…]
(a) A simplex equiangular tight frame
(b) Gradient directions w.r.t. $\bm{w}$
(c) Gradient directions w.r.t. $\bm{h}$
Figure 3. Gradient analysis based on the neural collapse theory. (a) An illustration of a simplex ETF when $d=3$ and $K=4$. The “+” and “” with different colors refer to features and classifier vectors of different classes, respectively. (b) Gradient directions about a certain $\bm{h}$ (belongs to the 1-st class) w.r.t. all $\bm{w}_{i},i\in{1,2,3}$. (c) Gradient directions w.r.t. an $\bm{h}$ (belongs to the 1-st class).
[…]
@@ 3.1. Incompatibility between CE and BinaryKL
$\mathcal{L}_{\mathrm{BinaryKL}}=\tau^{2}\sum\limits_{i=1}^{B}\sum\limits_{k=1}^{K}\mathcal{KL}\Big{(}$ $\left[\sigma\left(z_{i,k}^{\mathcal{T}}/\tau\right),1-\sigma\left(z_{i,k}^{\mathcal{T}}/\tau\right)\right]$
[…]
@@
As shown in Figures 1(b) and 2, although the BinaryKL loss performs well alone, it is incompatible with the CE loss. The student model can achieve descent performance by training with either the CE loss or the BinaryKL loss separately, but the performance will be degraded severely when we combine them as follows:
(2) $\mathcal{L}_{\mathrm{overall}}=\mathcal{L}_{\mathrm{CE}}+\alpha\mathcal{L}_{\mathrm{BinaryKL}}.$
@@
Figure 4. An illustration of Dual-Head Knowledge Distillation (DHKD). DHKD decouples the original linear classifier into a duo. We can evade conflicts between two losses by introducing an Adaptive Auxiliary Classifier customized according to the models’ architectures. The right side shows the gradient alignment method on CIFAR-100. When the angle between two gradients is larger than $90^{\circ}$, we will project the gradient of the BinaryKL loss to the orthogonal direction of the gradient of the CE loss.
[…]
3.2. Theoretical Analysis
@@
Papyan et al. (2020) revealed the neural collapse phenomenon, where the last-layer features converge to their within-class means, and the within-class means together with the classifier vectors collapse to the vertices of a simplex equiangular tight frame (ETF) at the terminal phase of training. According to Papyan et al. (2020), all vectors in a simplex ETF have an equal $\ell_{2}$ norm and the same pairwise angle. An illustration of a simplex ETF is shown in Figure 3(c). The neural collapse phenomenon (Papyan et al., 2020) can be characterized by four manifestations: (1) the variability of the last-layer features in the same class collapses to zero; (2) the class means of different classes’ last-layer features converge to a simplex ETF; (3) the linear classifier vectors collapse to the class means; (4) the task of predicting for classification can be simplified into finding out the nearest class center of the last-layer feature. The detailed definition of a simplex ETF and the features of the neural collapse phenomenon can be found in Appendix B.1.
@@
(3) $\frac{\partial\mathcal{L}_{\mathrm{overall}}}{\partial\bm{w}_{k}}=-(\bm{F}_{\bm{w}\mathrm{\text{-}pull}}^{\mathrm{CE}}+\alpha\bm{F}_{\bm{w}\mathrm{\text{-}pull}}^{\mathrm{BinaryKL}})-(\bm{F}_{\bm{w}\mathrm{\text{-}push}}^{\mathrm{CE}}+\alpha\bm{F}_{\bm{w}\mathrm{\text{-}push}}^{\mathrm{BinaryKL}}),$
@@
$\bm{F}_{\bm{w}\mathrm{\text{-}pull}}^{\mathrm{CE}}=\sum\nolimits_{i=1}^{n_{k}}(1-p_{k}(\bm{h}^{\mathcal{S}}_{k,i}))\bm{h}^{\mathcal{S}}_{k,i},$
$\bm{F}_{\bm{w}\mathrm{\text{-}pull}}^{\mathrm{BinaryKL}}=\tau\sum\nolimits_{i=1}^{n_{k}}(q_{k}(\bm{h}^{\mathcal{T}}_{k,i})-q_{k}(\bm{h}^{\mathcal{S}}_{k,i}))\bm{h}^{\mathcal{S}}_{k,i},$
$\bm{F}_{\bm{w}\mathrm{\text{-}push}}^{\mathrm{CE}}=-\sum\nolimits_{k^{\prime}\neq k}^{K}\sum\nolimits_{j=1}^{n_{k^{\prime}}}p_{k}(\bm{h}^{\mathcal{S}}_{k^{\prime},i})\bm{h}^{\mathcal{S}}_{k^{\prime},j},$
(4) $\bm{F}_{\bm{w}\mathrm{\text{-}push}}^{\mathrm{BinaryKL}}=-\tau\sum\nolimits_{k^{\prime}\neq k}^{K}\sum\nolimits_{j=1}^{n_{k^{\prime}}}(q_{k}(\bm{h}^{\mathcal{S}}_{k,i})-q_{k}(\bm{h}^{\mathcal{T}}_{k,i}))\bm{h}^{\mathcal{S}}_{k^{\prime},j}.$
@@
(5) $\frac{\partial\mathcal{L}_{\mathrm{overall}}}{\partial\bm{h}}=-(\bm{F}_{\bm{h}\mathrm{\text{-}pull}}^{\mathrm{CE}}+\alpha\bm{F}_{\bm{h}\mathrm{\text{-}pull}}^{\mathrm{BinaryKL}})-(\bm{F}_{\bm{h}\mathrm{\text{-}push}}^{\mathrm{CE}}+\alpha\bm{F}_{\bm{h}\mathrm{\text{-}push}}^{\mathrm{BinaryKL}}),$
@@
$\bm{F}_{\bm{h}\mathrm{\text{-}pull}}^{\mathrm{CE}}=(1-p_{c}(\bm{h}^{\mathcal{S}}))\bm{w}^{\mathcal{S}}_{c},$
$\bm{F}_{\bm{h}\mathrm{\text{-}pull}}^{\mathrm{BinaryKL}}=\tau(q_{c}(\bm{h}^{\mathcal{T}})-q_{c}(\bm{h}^{\mathcal{S}}))\bm{w}^{\mathcal{S}}_{c},$
$\bm{F}_{\bm{h}\mathrm{\text{-}push}}^{\mathrm{CE}}=-\sum\nolimits_{k\neq c}^{K}p_{k}(\bm{h}^{\mathcal{S}})\bm{w}^{\mathcal{S}}_{k},$
(6) $\bm{F}_{\bm{h}\mathrm{\text{-}push}}^{\mathrm{BinaryKL}}=-\tau\sum\nolimits_{k\neq c}^{K}(q_{k}(\bm{h}^{\mathcal{S}})-q_{k}(\bm{h}^{\mathcal{T}}))\bm{w}^{\mathcal{S}}_{k}.$
@@
Table 1. Results on the CIFAR-100 validation. Teachers and students are in the same architectures. The best performance of a single method is highlighted with a star , and the best logits-based performance is highlighted in bold. The combination of DHKD and ReviewKD always achieves the best performance and is highlighted with gray backgrounds.
[…]
. Teacher resnet56 resnet110 resnet32$\times$4 WRN-40-2 WRN-40-2 VGG13 72.34 74.31 79.42 75.61 75.61 74.64 Student resnet20 resnet32 resnet8$\times$4 WRN-16-2 WRN-40-1 VGG8 69.06 71.14 72.50 73.26 71.98 70.36 features FitNet 69.21 71.06 73.50 73.58 72.24 71.02 RKD 69.61 71.82 71.90 73.35 72.22 71.48 CRD 71.16 73.48 75.51 75.48 74.14 73.94 OFD 70.98 73.23 74.95 75.24 74.33 73.95 ReviewKD 71.89 73.89 75.63 76.12 75.09 74.84 SimKD 71.02 73.89 78.04 75.48 75.21 74.83 CAT-KD 71.62 73.62 76.91 75.60 74.82 74.65 logits KD 70.66 73.08 73.33 74.92 73.54 72.98 DKD $\textbf{71.97}^{*}$ $\textbf{74.11}^{*}$ 76.32 76.24 74.81 74.68 DHKD 71.19 73.92 76.54 $\textbf{76.36}^{*}$ $\textbf{75.25}^{*}$ $\textbf{74.84}^{*}$ DHKD + ReviewKD 73.14 75.21 78.29 77.97 76.56 76.27
3.3. Our Proposed Method